Suppose a country's real GDP is $14 trillion and that population is 200 million 

Instructions: Enter your answers as whole numbers 


a. What is this country's real GDP per capita? Suppose that during the next 10 years, real GDP grows by half and population triples in the country. 


b. After 10 years have passed, what will be this country's real GDP per capita?
Answer #1
Answer #2

SOLUTION :


1 trillion = 10^6 million.


a.


Real GDP per capita 

= Real GDP / Population 

= 14*10^6 million $  /  200 million

= 70000 ($) (ANSWER)


b.


After 10 years :


New real GDP = 14*10^6 * 1.5 = 21*10^6 million $ 

New population = 3*200 = 600 million


So,


Real GDP per capita after 10 years

= New real GDP / New population

= 21*10^6 million $ / 600 million.

= 35000 ($) (ANSWER).

answered by: Tulsiram Garg